We had a telescope driver failure at the beginning of the night. We also had several other technical issues: a shutter hardware error and an alert "Component not ready for exposure: DECAL."
After the driver failure was rectified, we had a few exposures with very large pointing offsets. Afterwards, the pointings normalized, but about an hour later we had another brief period where the offsets grew large (~140") in dec (RA was mostly ok). There was nothing else out of the ordinary otherwise during this period (ie, we were not making large slews).
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-03-28 | 8.5 | 0.0 | 1.5 | 0.0 | 0.0 | 10 |
Links
[1] mailto:djschlegel@lbl.gov
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/2245/delete?destination=print/content/enr-ctio-blanco-4-m-2015-03-28
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/2245/edit?destination=print/content/enr-ctio-blanco-4-m-2015-03-28
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/7489?destination=print/content/enr-ctio-blanco-4-m-2015-03-28
Water cooler in kitchen is empty.
We also observed a Target of Opportunity for NEOWISE (prop id: 2015A-0059).